# Label Distribution Learning using the Squared Neural Family on the Probability Simplex

### v1

@@ Appendix A Theorem Proofs
Appendix B Label Distribution Prediction Results
Tables 5-9 show the label distribution prediction performance measured by Chebyshev distance, Clark distance, Canberra metric, Kullback-Leibler divergence and Intersection, respectively, as the averaged scores on ten-fold cross-validation. On each dataset, the best performer and runner-up are respectively highlighted by boldface and underline. We perform paired t tests between the proposed SNEFY-LDL and every baseline method, and use symbols $\bullet$, $\circ$ and $\spadesuit$ to respectively denote that SNEFY-LDL is superior, inferior and equivalent to the baseline at 0.05 significance level. The last column $\Delta$ gives the performance difference between SNEFY-LDL and the best baseline. The scores in Tables 5-9 consistently confirm that SNEFY-LDL is very competitive compared with the state-of-the-art baselines.
Dataset PT-SVM AA-KNN CPNN SA-BFGS LDLF DF-LDL LDL-SCL LDL-LRR SNEFY-LDL $\Delta$
Movie 0.1183 $\bullet$ 0.1193 $\bullet$ 0.1330 $\bullet$ 0.1163 $\spadesuit$ 0.1668 $\bullet$ 0.1135 $\circ$ 0.1444 $\bullet$ 0.1154 $\circ$ 0.1172 +0.0037
Natural_Scene 0.3104 $\spadesuit$ 0.3130 $\spadesuit$ 0.3563 $\bullet$ 0.3194 $\spadesuit$ 0.3023 $\spadesuit$ 0.3077 $\spadesuit$ 0.3603 $\bullet$ 0.3046 $\spadesuit$ 0.3135 +0.0112
SBU_3DFE 0.1300 $\bullet$ 0.1277 $\bullet$ 0.1394 $\bullet$ 0.1152 $\bullet$ 0.1197 $\bullet$ 0.1147 $\bullet$ 0.1303 $\bullet$ 0.1268 $\bullet$ 0.1110 -0.0037
SJAFFE 0.1158 $\bullet$ 0.0975 $\spadesuit$ 0.1202 $\bullet$ 0.0921 $\spadesuit$ 0.0862 $\circ$ 0.0810 $\circ$ 0.0868 $\circ$ 0.0831 $\circ$ 0.0997 +0.0188
Yeast_alpha 0.0140 $\bullet$ 0.0144 $\bullet$ 0.0140 $\bullet$ 0.0134 $\circ$ 0.0138 $\spadesuit$ 0.0134 $\circ$ 0.0135 $\circ$ 0.0134 $\circ$ 0.0137 +0.0003
Yeast_cdc 0.0168 $\spadesuit$ 0.0172 $\bullet$ 0.0169 $\bullet$ 0.0162 $\circ$ 0.0163 $\circ$ 0.0162 $\circ$ 0.0164 $\circ$ 0.0162 $\circ$ 0.0167 +0.0005
Yeast_cold 0.0529 $\bullet$ 0.0542 $\bullet$ 0.0538 $\bullet$ 0.0512 $\spadesuit$ 0.0535 $\bullet$ 0.0514 $\spadesuit$ 0.0543 $\bullet$ 0.0512 $\spadesuit$ 0.0518 +0.0006
Yeast_diau 0.0386 $\spadesuit$ 0.0384 $\spadesuit$ 0.0407 $\bullet$ 0.0369 $\circ$ 0.0376 $\spadesuit$ 0.0369 $\circ$ 0.0405 $\bullet$ 0.0370 $\circ$ 0.0382 +0.0013
Yeast_dtt 0.0381 $\bullet$ 0.0385 $\bullet$ 0.0370 $\spadesuit$ 0.0360 $\circ$ 0.0374 $\spadesuit$ 0.0365 $\spadesuit$ 0.0393 $\spadesuit$ 0.0360 $\circ$ 0.0369 +0.0009
Yeast_elu 0.0171 $\bullet$ 0.0173 $\bullet$ 0.0169 $\spadesuit$ 0.0163 $\circ$ 0.0163 $\circ$ 0.0162 $\circ$ 0.0164 $\circ$ 0.0163 $\circ$ 0.0168 +0.0005
Yeast_heat 0.0438 $\bullet$ 0.0440 $\bullet$ 0.0433 $\bullet$ 0.0422 $\spadesuit$ 0.0425 $\spadesuit$ 0.0420 $\circ$ 0.0435 $\spadesuit$ 0.0422 $\spadesuit$ 0.0425 +0.0006
Yeast_spo 0.0589 $\bullet$ 0.0629 $\bullet$ 0.0593 $\bullet$ 0.0584 $\spadesuit$ 0.0612 $\bullet$ 0.0603 $\bullet$ 0.0592 $\bullet$ 0.0583 $\spadesuit$ 0.0583 -0.0000
Yeast_spo5 0.0923 $\bullet$ 0.0952 $\bullet$ 0.0920 $\bullet$ 0.0913 $\spadesuit$ 0.0998 $\bullet$ 0.0973 $\bullet$ 0.1083 $\bullet$ 0.0912 $\spadesuit$ 0.0910 -0.0002
Yeast_spoem 0.0884 $\bullet$ 0.0907 $\bullet$ 0.0874 $\spadesuit$ 0.0868 $\spadesuit$ 0.0932 $\bullet$ 0.1043 $\bullet$ 0.2315 $\bullet$ 0.0865 $\spadesuit$ 0.0865 -0.0000
Table 5: The comparison of label distribution prediction performance measured by Chebyshev $\downarrow$.
Dataset PT-SVM AA-KNN CPNN SA-BFGS LDLF DF-LDL LDL-SCL LDL-LRR SNEFY-LDL $\Delta$
Movie 0.5293 $\circ$ 0.5312 $\spadesuit$ 0.5636 $\bullet$ 0.5251 $\circ$ 0.7024 $\bullet$ 0.5107 $\circ$ 0.6230 $\bullet$ 0.5214 $\circ$ 0.5370 +0.0263
Natural_Scene 2.4441 $\circ$ 1.9084 $\circ$ 2.4878 $\bullet$ 2.4234 $\circ$ 2.4313 $\circ$ 2.4130 $\circ$ 2.3683 $\circ$ 2.4399 $\circ$ 2.4568 +0.5484
SBU_3DFE 0.4067 $\bullet$ 0.4015 $\bullet$ 0.4151 $\bullet$ 0.3710 $\bullet$ 0.3770 $\bullet$ 0.3624 $\bullet$ 0.4025 $\bullet$ 0.3975 $\bullet$ 0.3429 -0.0195
SJAFFE 0.4167 $\bullet$ 0.3477 $\spadesuit$ 0.4293 $\bullet$ 0.3589 $\spadesuit$ 0.3350 $\circ$ 0.3031 $\circ$ 0.3222 $\circ$ 0.3149 $\circ$ 0.3613 +0.0582
Yeast_alpha 0.2218 $\bullet$ 0.2261 $\bullet$ 0.2239 $\bullet$ 0.2099 $\circ$ 0.2166 $\spadesuit$ 0.2094 $\circ$ 0.2122 $\circ$ 0.2099 $\circ$ 0.2182 +0.0088
Yeast_cdc 0.2255 $\bullet$ 0.2317 $\bullet$ 0.2243 $\bullet$ 0.2158 $\circ$ 0.2173 $\circ$ 0.2157 $\circ$ 0.2179 $\circ$ 0.2159 $\circ$ 0.2219 +0.0062
Yeast_cold 0.1443 $\bullet$ 0.1477 $\bullet$ 0.1458 $\bullet$ 0.1396 $\spadesuit$ 0.1456 $\bullet$ 0.1405 $\spadesuit$ 0.1479 $\bullet$ 0.1396 $\spadesuit$ 0.1411 +0.0015
Yeast_diau 0.2096 $\bullet$ 0.2080 $\spadesuit$ 0.2178 $\bullet$ 0.2007 $\circ$ 0.2032 $\spadesuit$ 0.2001 $\circ$ 0.2213 $\bullet$ 0.2007 $\circ$ 0.2070 +0.0068
Yeast_dtt 0.1041 $\bullet$ 0.1042 $\bullet$ 0.1008 $\spadesuit$ 0.0982 $\circ$ 0.1017 $\spadesuit$ 0.0994 $\spadesuit$ 0.1065 $\spadesuit$ 0.0982 $\circ$ 0.1006 +0.0025
Yeast_elu 0.2114 $\bullet$ 0.2131 $\bullet$ 0.2073 $\spadesuit$ 0.1990 $\circ$ 0.2003 $\circ$ 0.1985 $\circ$ 0.2014 $\circ$ 0.1990 $\circ$ 0.2077 +0.0092
Yeast_heat 0.1893 $\bullet$ 0.1908 $\bullet$ 0.1876 $\bullet$ 0.1826 $\spadesuit$ 0.1844 $\spadesuit$ 0.1815 $\circ$ 0.1882 $\bullet$ 0.1826 $\spadesuit$ 0.1839 +0.0024
Yeast_spo 0.2530 $\bullet$ 0.2665 $\bullet$ 0.2533 $\bullet$ 0.2498 $\spadesuit$ 0.2594 $\bullet$ 0.2552 $\bullet$ 0.2530 $\bullet$ 0.2497 $\spadesuit$ 0.2494 -0.0003
Yeast_spo5 0.1861 $\bullet$ 0.1915 $\bullet$ 0.1857 $\bullet$ 0.1840 $\spadesuit$ 0.1999 $\bullet$ 0.1951 $\bullet$ 0.2145 $\bullet$ 0.1838 $\spadesuit$ 0.1833 -0.0005
Yeast_spoem 0.1313 $\bullet$ 0.1350 $\bullet$ 0.1299 $\spadesuit$ 0.1292 $\spadesuit$ 0.1392 $\bullet$ 0.1541 $\bullet$ 0.3628 $\bullet$ 0.1287 $\spadesuit$ 0.1288 +0.0001

### v2

@@ Appendix A Theorem Proofs
Appendix B Experimental Details
B.1 Conformal Prediction
We select the Natural Scene and SBU_3DFE datasets for the conformal prediction experiments, as they have a relatively large number of labels that give more freedom in the change of label distribution compositions to warrant more uncertainties in label distribution predictions, and they also have enough samples to do the training/calibration/test set split. According to the ratio of 50%/25%/25%, we randomly split each dataset into the training, calibration and test sets. Here, we denote the training, calibration and test sets as $\mathcal{S}^{\text{tr}}=\{(\bm{x}^{\text{tr}}_{i},\bm{\ell}_{\bm{x}^{\text{tr}}_{i}}):i=1,\cdots,N_{\text{tr}}\}$, $\mathcal{S}^{\text{cal}}=\{(\bm{x}^{\text{cal}}_{i},\bm{\ell}_{\bm{x}^{\text{cal}}_{i}}):i=1,\cdots,N_{\text{cal}}\}$ and $\mathcal{S}^{\text{te}}=\{(\bm{x}^{\text{te}}_{i},\bm{\ell}_{\bm{x}^{\text{te}}_{i}}):i=1,\cdots,N_{\text{te}}\}$ respectively, with $N_{\text{tr}}$, $N_{\text{cal}}$ and $N_{\text{te}}$ denoting the number of samples in the training, calibration and test sets respectively. For each split, a SNEFY-LDL is first trained on the training set $\mathcal{S}^{\text{tr}}$. The trained SNEFY-LDL is then leveraged to construct a 90% level confidence interval for each label’s composition ratio with SNEFY-LDL’s closed-form conditional mean and variance. The confidence intervals constructed by the trained SNEFY-LDL model are then calibrated with the calibration set. The adaptivity of the calibrated confidence intervals are finally evaluated on the test set, as an indicator of SNEFY-LDL’s ability in quantifying the prediction uncertainty of each label’s composition ratio. On the test set, the adaptivity of the calibrated confidence intervals $\mathcal{C}(\bm{x},\bm{\ell}^{y_{r}})$ in Eq. (19) for the $r$th label’s composition ratio $\bm{\ell}^{y_{r}}$ is measured by the Feature-Stratified Coverage (FSC) metric [Angelopoulos and Bates, 2021]. For calculating the FSC metric, we first bin the first feature of $\bm{x}$ for all $\bm{x}\in\mathcal{S}_{\text{te}}$ into a number of categories, $1,\cdots,G$, then categorize test samples in $\mathcal{S}_{\text{te}}$ into different groups $\{\mathcal{S}^{\text{te}}_{g}:g=1,\cdots,G\}$ according to the first feature’s category values. Here, $G$ is termed as bin size or group number. The FSC metric for the confidence intervals of the $r$th label’s composition ratio $\bm{\ell}^{y_{r}}$ is evaluated as the minimal coverage rate among the groups $\{\mathcal{S}^{\text{te}}_{g}:g=1,\cdots,G\}$:
$\text{FSC}(\bm{\ell}^{y_{r}}):=\min_{g\in\{1,\cdots,G\}}\frac{1}{|\mathcal{S}_{g}^{\text{te}}|}\sum_{\bm{x}\in\mathcal{S}_{g}^{\text{te}}}\mathbbm{1}\left\{\bm{\ell}_{\bm{x}}^{y_{r}}\in\mathcal{C}(\bm{x},\bm{\ell}^{y_{r}})\right\}.$ (20)
The detailed procedure of conformal prediction with SNEFY-LDL is described by Algorithm A1.
Algorithm A1 Conformal Prediction with SNEFY-LDL
Input: Training set $\mathcal{S}^{\text{tr}}=\{(\bm{x}^{\text{tr}}_{i},\bm{\ell}_{\bm{x}^{\text{tr}}_{i}}):i=1,\cdots,N_{\text{tr}}\}$, calibration set $\mathcal{S}^{\text{cal}}=\{(\bm{x}^{\text{cal}}_{i},\bm{\ell}_{\bm{x}^{\text{cal}}_{i}}):i=1,\cdots,N_{\text{cal}}\}$ and test set $\mathcal{S}^{\text{te}}=\{(\bm{x}^{\text{te}}_{i},\bm{\ell}_{\bm{x}^{\text{te}}_{i}}):i=1,\cdots,N_{\text{te}}\}$, as well as the given label set $\mathcal{Y}=\{y_{1},y_{2},\cdots,y_{L}\}$.
Parameter: $0.11/k^{2}=$ 0.1 for constructing 90% level confidence intervals.
Output: FSC scores on the test set.
1:  Train a SNEFY-LDL model on the training set $\mathcal{S}^{\text{tr}}$ with Algorithm 1;
2:  for each sample $\bm{x}\in\mathcal{S}^{\text{cal}}$ do
3:     for each label $y_{r}\in\mathcal{Y}$ do
4:        Calculate the calibration score $s(\bm{x},\bm{\ell}^{y_{r}}_{\bm{x}})$ with Eq.(18) and the trained SNEFY-LDL model;
5:     end for
6:  end for
7:  for each label $y_{r}\in\mathcal{Y}$ do
8:     Calculate the $\lceil(1-1/k^{2})(N_{\text{cal}}+1)\rceil/N_{\text{cal}}$ quantile as $\hat{q}_{y_{r}}$ among the $N_{\text{cal}}$ calibration scores $s(\bm{x},\bm{\ell}^{y_{r}}_{\bm{x}})$ with $\bm{x}\in\mathcal{S}^{\text{cal}}$;
9:  end for
10:  for each sample $\bm{x}\in\mathcal{S}^{\text{te}}$ do
11:     for each label $y_{r}\in\mathcal{Y}$ do
12:        Construct the calibrated confidence interval $\mathcal{C}(\bm{x},\bm{\ell}^{y_{r}})$ with Eq. (19) and the calculated quantile $\hat{q}_{y_{r}}$;
13:     end for
14:  end for
15:  for each label $y_{r}\in\mathcal{Y}$ do
16:     Evaluate the $\text{FSC}(\bm{\ell}^{y_{r}})$ score with Eq. (20) and the $N_{\text{te}}$ confidence intervals $\mathcal{C}(\bm{x},\bm{\ell}^{y_{r}})$ with $\bm{x}\in\mathcal{S}^{\text{te}}$;
17:  end for
18:  return the evaluated $\text{FSC}(\bm{\ell}^{y_{r}})$ scores for all $y_{r}\in\mathcal{Y}$.
B.2 Active Learning
The Motive and Natural Scene datasets are selected for active learning, as they are relatively sensitive to the label sparsity issue, more suitable to benchmark the performance change with informative samples labeled and augmented to the training data. We randomly split the two datasets into the training and test sets according to the ratio of 90%/10% for ten times. We denote the training set as $\mathcal{S}^{\text{tr}}=\{(\bm{x}^{\text{tr}}_{i},\bm{\ell}_{\bm{x}^{\text{tr}}_{i}}):i=1,\cdots,N_{\text{tr}}\}$ and test set as $\mathcal{S}^{\text{te}}=\{(\bm{x}^{\text{te}}_{i},\bm{\ell}_{\bm{x}^{\text{te}}_{i}}):i=1,\cdots,N_{\text{te}}\}$, where $N_{\text{tr}}$ and $N_{\text{te}}$ are respectively the number of samples in the training and test sets. For each training-test set split, we randomly select 400 labeled samples from the training set $\mathcal{S}^{\text{tr}}$ to form the initial labeled pool and take the remaining samples in $\mathcal{S}^{\text{tr}}$ as unlabeled samples. We first train a SNEFY-LDL model with the initial labeled pool. To achieve active learning with the trained SNEFY-LDL model, we first evaluate the differential entropy $H(\bm{x};\bm{V},\bm{\Theta})$ for each unlabeled sample $\bm{x}$ as
$H(\bm{x};\bm{V},\bm{\Theta})=-\int_{\Delta^{L-1}}\left\{\log\frac{\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})}{d\bm{\ell}}\right\}\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta}),$ (21)
then select 100 most informative unlabeled samples with the largest differential entropy values. After querying the labels of the 100 selected samples, we augment them into the initial labeled pool, and re-train another SNEFY-LDL model with the augmented labeled pool. With the re-trained SNEFY-LDL model, we can predict the label distribution vectors for samples in the test set given their feature vectors $\bm{x}$ by directly using the closed-form conditional mean in Eq.(10). As a criterion of active learning, the label distribution prediction performance on the test set is evaluated using the six metrics [Geng, 2016]: Chebyshev distance (Cheby), Clark distance (Clark), Canberra metric (Canb), Kullback-Leibler divergence (KL), Cosine coefficient (Cos) and Intersection (Inter). Given the ground-truth and predicted label distribution vectors as $\bm{\ell}$ and $\hat{\bm{\ell}}$ respectively, the evaluation metrics are defined as follows:
$\text{Cheby}(\bm{\ell},\hat{\bm{\ell}})=\|\bm{\ell}-\hat{\bm{\ell}}\|_{\infty}\downarrow,\quad\text{Clark}(\bm{\ell},\hat{\bm{\ell}})=\Big{\|}\frac{\bm{\ell}-\hat{\bm{\ell}}}{\bm{\ell}+\hat{\bm{\ell}}}\Big{\|}_{2}\downarrow,\quad\text{Canb}(\bm{\ell},\hat{\bm{\ell}})=\Big{\|}\frac{\bm{\ell}-\hat{\bm{\ell}}}{\bm{\ell}+\hat{\bm{\ell}}}\Big{\|}_{1}\downarrow,$ (22)
$\text{KL}(\bm{\ell},\hat{\bm{\ell}})=\sum_{y\in\mathcal{Y}}\bm{\ell}^{y}\log\frac{\bm{\ell}^{y}}{\hat{\bm{\ell}}^{y}}\downarrow,\;\;\;\text{Cos}(\bm{\ell},\hat{\bm{\ell}})=\frac{\bm{\ell}^{\top}\hat{\bm{\ell}}}{\|\bm{\ell}\|_{2}\|\hat{\bm{\ell}}\|_{2}}\uparrow,\quad\;\;\text{Inter}(\bm{\ell},\hat{\bm{\ell}})=\sum_{y\in\mathcal{Y}}\min(\bm{\ell}^{y},\hat{\bm{\ell}}^{y})\uparrow.$
However, the differential entropy $H(\bm{x};\bm{V},\bm{\Theta})$ in Eq. (21) cannot be computed in a closed form. To overcome this difficulty, we adopt importance sampling [Kloek and Van Dijk, 1978] to approximately estimate the differential entropy values, where the uniform distribution is selected as the proposal distribution. The detailed procedure is provided in Algorithm A2.
Algorithm A2 Differential Entropy Estimation with Importance Sampling
Input: The feature vector of an unlabeled sample $\bm{x}$ and the conditional distribution $\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})$ modeled by the trained SNEFY-LDL model, and the label set size $L$.
Parameter: The number of sampling iterations $N_{\text{iter}}=1,000$.
Output: The estimated differential entropy for the unlabeled sample $\bm{x}$, $\hat{H}(\bm{x};\bm{V},\bm{\Theta})$.
1:  for each iteration $i\in\{1,\cdots,N_{\text{iter}}\}$ do
2:     Sample a label distribution vector ~$\tilde{\bm{\ell}}^{(i)}$ from the uniform distribution over the simplex $\Delta^{L-1}$ with probability density $q(\tilde{\bm{\ell}}^{(i)})=(L-1)!$;
3:     Evaluate the probability density value $\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})/d\bm{\ell}$ at =~$\bm{\ell}=\tilde{\bm{\ell}}^{(i)}$ as $p(\tilde{\bm{\ell}}^{(i)}|\bm{x};\bm{V},\bm{\Theta})$ with Eq. (6) and the trained SNEFY-LDL model;
4:  end for
5:  Calculate the approximated differential entropy $\hat{H}(\bm{x};\bm{V},\bm{\Theta})=-\frac{1}{N_{\text{iter}}}\sum_{i=1}^{N_{\text{iter}}}\frac{p(\tilde{\bm{\ell}}^{(i)}|\bm{x};\bm{V},\bm{\Theta})}{q(\tilde{\bm{\ell}}^{(i)})}\log p(\tilde{\bm{\ell}}^{(i)}|\bm{x};\bm{V},\bm{\Theta})$;
6:  return the estimated differential entropy $\hat{H}(\bm{x};\bm{V},\bm{\Theta})$ for the given unlabeled sample $\bm{x}$.
The detailed procedure for active learning with SNEFY-LDL is described by Algorithm A3.
B.3 Ensemble Learning
The SBU_3DFE and SJAFFE datasets are selected for the ensemble learning experiments, as they have a relatively small number of features, more efficient to train a number of base learners and do ensemble prediction. We randomly split the two datasets into the training and test sets according to the ratio of 90%/10% for ten times. We denote the training set as $\mathcal{S}^{\text{tr}}=\{(\bm{x}^{\text{tr}}_{i},\bm{\ell}_{\bm{x}^{\text{tr}}_{i}}):i=1,\cdots,N_{\text{tr}}\}$ and test set as $\mathcal{S}^{\text{te}}=\{(\bm{x}^{\text{te}}_{i},\bm{\ell}_{\bm{x}^{\text{te}}_{i}}):i=1,\cdots,N_{\text{te}}\}$, where $N_{\text{tr}}$ and $N_{\text{te}}$ are respectively the number of samples in the training and test sets. For each training-test set split, we randomly select 50 samples from the training set for 25 rounds, train 25 base LDL learners with the selected samples, and evaluate the label distribution prediction performance of the ensembled LDL model on the test set. Four competitive LDL algorithms are employed to train base learners: SA-BFGS [Geng, 2016], DF-LDL [González et al., 2021a], LDL-SCL [Jia et al., 2021] and LDL-LRR [Jia et al., 2023a]. To achieve ensemble learning with SNEFY-LDL, we first train a SNEFY-LDL model with the training set $\mathcal{S}^{\text{tr}}$, and then do the ensemble prediction by weighting the base learners according to the SNEFY-LDL conditional probability densities measured at their predictions given each test sample $\bm{x}\in\mathcal{S}^{\text{te}}$. The detailed procedure is shown in Algorithm A4.
Algorithm A3 Active Learning with SNEFY-LDL
Input: Training set $\mathcal{S}^{\text{tr}}=\{(\bm{x}^{\text{tr}}_{i},\bm{\ell}_{\bm{x}^{\text{tr}}_{i}}):i=1,\cdots,N_{\text{tr}}\}$ and test set $\mathcal{S}^{\text{te}}=\{(\bm{x}^{\text{te}}_{i},\bm{\ell}_{\bm{x}^{\text{te}}_{i}}):i=1,\cdots,N_{\text{te}}\}$.
Parameters: The size of the initial labeled pool $N_{\text{initial}}=400$ and the number of queried samples $N_{\text{query}}=100$.
Output: The label distribution prediction performance scores measured by Cheby, Clark, Canb, KL, Cos and Inter on the test set $\mathcal{S}^{\text{te}}$.
1:  Randomly select $N_{\text{initial}}$ samples from the training set $\mathcal{S}^{\text{tr}}$ to form the labeled pool $\mathcal{S}^{\text{tr}}_{\text{label}}$ and use the remaining samples to form the unlabeled pool $\mathcal{S}^{\text{tr}}_{\text{unlabel}}$;
2:  Train a SNEFY-LDL model with the initial labeled pool;
3:  for each sample $\bm{x}\in\mathcal{S}^{\text{tr}}_{\text{unlabel}}$ do
4:     Estimate the differential entropy value $\hat{H}(\bm{x};\bm{V},\bm{\Theta})$ for $\bm{x}$ with Algorithm A2 and the trained SNEFY-LDL model;
5:  end for
6:  Select $N_{\text{query}}$ samples from $\mathcal{S}^{\text{tr}}_{\text{unlabel}}$ with the top-$N_{\text{query}}$ differential entropy values and query their label distributions;
7:  Augment the $N_{\text{query}}$ queried samples into the labeled pool $\mathcal{S}^{\text{tr}}_{\text{label}}$;
8:  Re-train another SNEFY-LDL model with the augmented labeled pool $\mathcal{S}^{\text{tr}}_{\text{label}}$;
9:  Evaluate the label distribution prediction performance of the re-trained SNEFY-LDL model on the test set $\mathcal{S}^{\text{te}}$ (as an average over all test samples) with the metrics of Cheby, Clark, Canb, KL, Cos and Inter defined in Eq. (22);